# iBERT: Interpretable Style Embeddings via Sense Decomposition

### v1

@@ Abstract
We present iBERT (interpretable-BERT), an encoder to produce inherently interpretable and controllable embeddings - designed to modularize and expose the discriminative cues present in language, such as stylistic and semantic structure. Each input token is represented as a sparse, non‑negative mixture over $k$ context‑independent sense vectors, which can be pooled into sentence embeddings or used directly at the token level. This enables modular control over representation, before any decoding or downstream use.
To demonstrate our model’s interpretability, we evaluate it on a suite of style‑focused tasks. On the STEL benchmark, it improves style representation effectiveness by $\sim$8 points over SBERT-style baselines, while maintaining competitive performance on authorship verification. Because each embedding is a structured composition of interpretable senses, we highlight how specific style attributes - such as emoji use, formality, or misspelling can be assigned to specific sense vectors. While our experiments center on style, iBERT is not limited to stylistic modeling. Its structural modularity is designed to interpretably decompose whichever discriminative signals are present in the data — enabling generalization even when supervision blends stylistic and semantic factors.
iBERT: Interpretable Style Embeddings via Sense Decomposition
Vishal Anand1, Milad Alshomary2, Kathleen McKeown2 1Microsoft, Washington, USA 2Columbia University, New York, USA Correspondence: vishal.anand@microsoft.com
@@ 1 Introduction
Though evaluated on stylistic tasks, iBERT is not a style model. It is a general-purpose encoder with decomposable representations—suitable for domains where interpretability and control are prerequisites rather than afterthoughts. We will make the code publicly available upon acceptance.
@@ 3 Method
We present iBERT, a multi-sense encoder that produces interpretable embeddings by design. iBERT can operate as a token-level encoder (pretrained as MLM) or as a sentence encoder (v1–v3) trained with contrastive supervision. Each token is encoded as a sparse mixture over static sense vectors, which are then pooled to form globally composable sentence embeddings. This design enables explicit analysis, control, and manipulation of embeddings.
@@ 3.3.2 Stage 1: Style Representaion Learning
To train iBERT on the task of style analysis, we use contrastive triplets $\langle s,s^{+},s^{-}\rangle$ targeting stylistic variation and create iBERT-v1-v3. Training data is drawn from StyleSynth / StyleDistance by Patel et al. (2025): content-controlled, synthetic style triplets; and from author-labeled triplets with topic control by Wegmann et al. (2022).
@@
To ensure architectural parity, we also train Vanilla-BERT and then Vanilla-SBERT baselines using the two-stage setup: a BERT-base model pretrained on the same 5% FineWeb slice, followed by contrastive fine-tuning on anchored triplets. This controls for pretraining variance and isolates the contribution of our multi-sense architecture.
@@
iBERT-v1–v3 has 171.4M parameters and Vanilla-SBERT has 149.7M. Inference is approximately 2% slower owing to sense composition. Tradeoffs are discussed in section 8.
@@ 4.1 Pretraining and Fine-tuning
We reuse iBERT-MLM and Vanilla-BERT models from Stage 0 (§3.3.1), trained from scratch on 5% of FineWeb. Sentence-level variants—iBERT-v1–v3 are fine-tuned on contrastive triplets to learn style representation using either:
  • StyleDistance - SD Patel et al. (2025): 50k synthetic triplets (SynthSTEL) with controlled lexical and syntactic variation. Median token-size is 19 (Appendix Table 7).
@@
Training follows the loss function of each dataset: for StyleDistance we use the InfoNCE objective as in Patel et al. (2025), while for Wegmann triplets we apply the margin-based contrastive loss used in Wegmann et al. (2022). In both cases, pooling weights $\pi_{\ell}(\tau)$ are shared between the anchor and positive to enforce consistent sense usage. Additional details are in Appendix A.4.
@@
To isolate the effect of our multi-sense architecture, we compare against contrastively trained versions of the same Vanilla-BERT encoder from Stage 0. We fine-tune this model to create:
  • Vanilla-BERTSD: trained on StyleDistance triplets with InfoNCE loss (as in Patel et al. (2025));
  • Vanilla-BERTWG: trained on Wegmann triplets with the margin-based triplet loss (as in Wegmann et al. (2022));
  • Vanilla-BERTWG+SD: trained on both datasets with InfoNCE loss (reflecting the strongest setup reported by Patel et al. (2025)).
Vanilla-BERTWG corresponds to the SBERT-style model of Wegmann et al. (2022), while Vanilla-BERTWG+SD mirrors the best-performing StyleDistance setup of Patel et al. (2025). All settings share the same backbone, initialization, and optimizer for a controlled comparison.
@@
Prior work used DistilRoBERTa-based encoders (e.g., Patel et al. (2025), Wegmann et al. (2022)). Since iBERT is trained from scratch, we also train a new BERT-base encoder using the same setup as iBERT-MLM. This serves as the backbone for Vanilla-BERTSD / WG / WG+SD, replacing DistilRoBERTa in their respective setups and enabling a fair performance comparison. Full training protocol is detailed in Appendix A.4.
@@ 4.2 Evaluation Benchmarks
  • STEL: multi-class classification over 40 style labels (StyleDistance) and 5 style classes (Wegmann), reported as weighted accuracy.
@@ 5.1 STEL Performance
Table 1: Direct-style: STEL and SoC reported on the 45 style groups (40 of StyleDistance + 5 of Wegmann STEL). Proxy-style: PAN reports average AUC% on PAN11/13/14/15 authorship verification tasks. The models are trained on corresponding Data triplets for 5 runs. Median token-lengths: 19 for SD, and 24 for WG (Appendix Table 7).
Table 1 reports style representation effectiveness (STEL) across training regimes Wegmann and Nguyen (2021). Across the board, iBERT models outperform their Vanilla counterparts, underscoring the utility of structured, interpretable pooling for capturing stylistic signals. Under SD-only training, iBERT-v3-1 achieves the highest STEL scores, yielding 38.3% (128 tokens) and 38.2% (512 tokens)—a substantial gain of +6.7 and +7.6 points over Vanilla. Close contenders include iBERT-v1 and iBERT-v3-10, both of which exhibit similarly strong performance, suggesting that soft, sense-weighted pooling is especially effective when guided by direct-style supervision.
@@
Even under joint supervision from WG+SD, where noisy author-level triplets may dilute sense-specific stylistic signal, iBERT variants maintain their edge. iBERT-v1 secures the best STEL at 128 tokens (38.6%), while iBERT-v3-1 leads at 512 tokens (36.4%), outperforming Vanilla by approximately 6 points in both cases.
Finally, in the WG-only setting, where labels reflect author identity rather than explicit styles, all models see degraded performance. Nevertheless, iBERT-v3 and iBERT-v1 is at par with Vanilla, signaling that even without style supervision, iBERT retains competitive structure-aware generalizability. These results affirm that iBERT’s compositional inductive bias enables robustness without compromising direct-style effectiveness.
@@ 5.2 SoC Performance
In contrast, iBERT-v2 underperforms across settings (e.g., 88.7% under SD-only), often lagging behind both v1 and Vanilla. We attribute this to its top-1 sense pooling: by forcing each sequence to align with a single dominant sense, iBERT-v2 severely restricts the number of senses that can specialize for the same data-points observed. This architectural bottleneck reduces expressive capacity, especially for fine-grained polarity distinctions that require distributed compositionality.
Under WG-only training—where supervision is provided weakly via author labels rather than direct stylistic polarity—all models see a dramatic drop in SoC performance. Vanilla, for example, falls from 91.5% (SD-only) to just 27.5%, while iBERT variants drop even further (e.g., 8.0% for v1). Despite this collapse, Vanilla appears to outperform iBERT in this setting. We hypothesize this is due to content leakage: Vanilla implicitly entangles semantic and stylistic attributes, using topic or lexical cues as proxies for style. While this can yield fragile SoC signal in weakly supervised settings, it sacrifices interpretability and modularity. In contrast, iBERT’s explicit disentanglement fails to recover polarity distinctions without true stylistic supervision—a tradeoff aligned with its core design goals.
@@ 5.3 Authorship Verification (PAN)
WG training dataset is a mixed dataset, containing author identity supervision with intertwined semantic and stylistic signals. Unlike Vanilla, which entangles all available cues, iBERT is architecturally designed to tease apart these differentials in an interpretable fashion. This enables it to maintain competitive performance even when the supervision is noisy or not explicitly style-specific—while still excelling on clean, style-focused tasks like STEL and SoC.
At 128 tokens, all iBERT variants match or slightly outperform Vanilla, with iBERT-v3-10 achieving the highest AUC (59.87%) for WG. This is notable not because iBERT ignores semantic information, but because it selectively organizes it in disentangled form—generalizing well even when stylistic cues are weakly defined.
At 512 tokens, Vanilla slightly outperforms iBERT (61.92% vs. 61.07% for v3), with the largest margins on PAN13 and PAN15, that are characterized by semantic drift and cross-topic variation. This suggests that Vanilla benefits more from extended context, due to its architecture’s tight coupling of semantic and stylistic features. In contrast, iBERT’s modular structure—designed to isolate stylistic axes—does not directly exploit semantic continuity. Still, the gap stays slim, highlighting the semantic signals acquired in MLM stage remain accessible even in disentangled representations.
@@ 5.4 Interpretability Analysis of iBERT
To validate these groupings, Table 3 reports the change in cosine distance between opposing style centroids before and after removing each sense contribution in final representation. When ablating a style-aligned sense, the separability between style polarities drops substantially (e.g., $66\%$ for $\ell=3$), confirming that key stylistic information are isolated within that sense. In contrast, most non-target styles exhibit low or negligible change in distance, indicating minimal collateral disruption: a hallmark of localized specialization.
@@
5.5 Visualizing Targeted Editing via Ablation
@@
For ablations on unrelated (non-target) senses, see Appendix A.6, which confirm that unrelated styles remain stable under targeted edits, affirming local modularity of learned sense vectors.
@@ 6 Discussion
Ablating a single sense (e.g., $\ell{=}1$ for emoji, or $\ell{=}4$ for nominalization) shifts the representation toward the negative style without disrupting unrelated styles. These transformations are visible both numerically (Table 3) and spatially (Fig. 2), confirming that iBERT learns axis-aligned subspaces that can be controlled with precision. This kind of intervention is infeasible in dense encoders like SBERT or Vanilla BERT.
@@ 7 Conclusion
iBERT matches or outperforms dense baselines across stylistic and semantic tasks like STEL, SoC, and PAN, while enabling: (a) disentangled encoding of style and semantics via sense-probing, (b) precise edits (e.g., emoji, formality), and (c) alignment of emergent senses with sociolinguistic axes.
@@
8 Limitations
@@ 9 Ethics Statement
iBERT is trained on public data, including synthetic edits (StyleDistance) and large-scale web corpora. As with all pretrained models, underlying biases in this data may influence the dimensions that emerge. We encourage careful evaluation before deployment in applications with societal or demographic impact.
@@ Appendix A Appendix
Figure 3: t-SNE projections of sentence embeddings for SynthSTEL. iBERT consistently separates contrastive style variants (e.g., misspelled sentences, emoji usage) better than Vanilla-SBERT, showing clearer margins and lower entanglement. Blue points represent positive samples and gray crosses represent negative samples. Despite iBERT-v2 being the most underperforming iBERT variant it still matches Vanilla’s performance, and the separation of positive and negative styles is cleaner.
(a) $\ell=0$ ablated that aligns with More/Less conjunctions
(b) $\ell=2$ ablated that aligns with Sarcasm / No sarcasm
Figure 4: All Upper Case/Proper Capitalization ($\ell$=1) maintains separation on ablating other styles. Colors as in Fig. 2- Red: original, Blue: edited, Gray: negatives.
@@
Table 5: Effect of aggregation sharpness $\tau$ on iBERT, trained on StyleDistance triplets (128 token size).
@@
Table 6: [iBERT-v3-10 (128-tokens)] Probing activations for style classes across all senses ($\ell=0$ to 77). The Best column indicates the latent sense with the highest activation for each style. Style names in this table reflect the original labels from the StyleDistance dataset. In the main text, shorter style names are used for presentation clarity.
@@
Variant Frozen Phase Unfrozen Phase Total
Time Loss / Epochs Time Loss / Epochs Time Epochs
128
iBERT 23h 1.6912 / 57 3d 6h 1.3365 / 85 4d 5h 142
BERT 2h 1.3607 / 5 1d 2h 1.2228 / 31 1d 5h 36
512
iBERT 2d 9h 1.4418 / 46 1d 7h 1.2941 / 16 3d 16h 62
BERT 14h 1.1876 / 6 3d 8h 1.0940 / 46 3d 22h 52
Table 8: Training runtimes, epochs, and final losses for iBERT-MLM and Vanilla-BERT (MLM) under 128- and 512-token limits. Epoch counts exclude early-stopping patience rounds.

### v2

# iBERT: Interpretable Embeddings via Sense Decomposition

@@ Abstract
We present iBERT (interpretable-BERT), an encoder to produce inherently interpretable and controllable embeddings - designed to modularize and expose the discriminative cues present in language, such as semantic or stylistic structure. Each input token is represented as a sparse, non‑negative mixture over $k$ context‑independent sense vectors, which can be pooled into sentence embeddings or used directly at the token level. This enables modular control over representation, before any decoding or downstream use.
To demonstrate our model’s interpretability, we evaluate it on a suite of style‑focused tasks. On the STEL benchmark, it improves style representation effectiveness by $\sim$8 points over SBERT-style baselines, while maintaining competitive performance on authorship verification. Because each embedding is a structured composition of interpretable senses, we highlight how specific style attributes get assigned to specific sense vectors. While our experiments center on style, iBERT is not limited to stylistic modeling. Its structural modularity is designed to interpretably decompose whichever discriminative signals are present in the data — enabling generalization even when supervision blends semantic or stylistic factors.
iBERT: Interpretable Embeddings via Sense Decomposition
Vishal Anand1, Milad Alshomary2, Kathleen McKeown2 1Microsoft, Washington, USA 2Columbia University, New York, USA Correspondence: vishal.anand@microsoft.com Website: https://iBERT.io
@@ 1 Introduction
Though evaluated on stylistic tasks, iBERT is not a style model. It is a general-purpose encoder with decomposable representations—suitable for domains where interpretability and control are prerequisites rather than afterthoughts.
@@ 3 Method
We present iBERT, a multi-sense encoder that produces interpretable embeddings by design. iBERT can operate as a token-level encoder (pretrained as MLM) or as a sentence encoder (v1–v3) trained with contrastive supervision. Each token is encoded as a sparse mixture over static sense vectors, which are then pooled to form globally composable sentence embeddings. This design enables explicit analysis, control, and manipulation of embeddings. In the paper, the usage of term BERT in a sentential context implies Sentence-BERT Reimers and Gurevych (2019).
@@ 3.3.2 Stage 1: Style Representaion Learning
To train iBERT on the task of style analysis, we use contrastive triplets $\langle s,s^{+},s^{-}\rangle$ targeting stylistic variation and create iBERT-v1-v3. Training data is drawn from StyleSynth by Patel et al. (2025): content-controlled, synthetic style triplets; and from author-labeled triplets with topic control by Wegmann et al. (2022).
@@
To ensure architectural parity, we also train BERT and then SBERT baselines using the two-stage setup: a BERT-base model pretrained on the same 5% FineWeb slice, followed by contrastive fine-tuning on anchored triplets. This controls for pretraining variance and isolates the contribution of our multi-sense architecture.
@@
iBERT-v1–v3 has 171.4M parameters and SBERT has 149.7M. Inference is approximately 2% slower owing to sense composition. Tradeoffs are discussed in Section Limitations.
@@ 4.1 Pretraining and Fine-tuning
We reuse iBERT-MLM and BERT models from Stage 0 (§3.3.1), trained from scratch on 5% of FineWeb. Sentence-level variants—iBERT-v1–v3 are fine-tuned on contrastive triplets to learn style representation using either:
  • StyleDistance - SD Patel et al. (2025): 50k synthetic triplets (StyleSynth) with controlled lexical and syntactic variation. Median token-size is 19 (Appendix Table 7).
@@
Training follows the loss function of each dataset: for StyleSynth we use the InfoNCE objective as in Patel et al. (2025), while for Wegmann triplets we apply the margin-based contrastive loss used in Wegmann et al. (2022). In both cases, pooling weights $\pi_{\ell}(\tau)$ are shared between the anchor and positive to enforce consistent sense usage. Additional details are in Appendix A.4.
@@
To isolate the effect of our multi-sense architecture, we compare against contrastively trained versions of the same BERT encoder from Stage 0. We fine-tune this model to create:
  • BERTSD: trained on StyleSynth triplets with InfoNCE loss (as in Patel et al. (2025));
  • BERTWG: trained on Wegmann triplets with the margin-based triplet loss (as in Wegmann et al. (2022));
  • BERTWG+SD: trained on both datasets with InfoNCE loss (reflecting the strongest setup reported by Patel et al. (2025)), which is listed as the prior state-of-the-art performing model called StyleDistance.
BERTWG corresponds to the SBERT-style model of Wegmann et al. (2022), while BERTWG+SD mirrors the best-performing StyleDistance setup of Patel et al. (2025). All settings share the same backbone, initialization, and optimizer for a controlled comparison.
@@
Prior work used DistilRoBERTa-based encoders (e.g., Patel et al. (2025), Wegmann et al. (2022)). Since iBERT is trained from scratch, we also train a new BERT-base encoder using the same setup as iBERT-MLM. This serves as the backbone for BERTSD / WG / WG+SD, replacing DistilRoBERTa in their respective setups and enabling a fair performance comparison. Full training protocol is detailed in Appendix A.4.
@@ 4.2 Evaluation Benchmarks
  • STEL: multi-class classification over 40 style labels (StyleSynth) and 5 style classes (Wegmann), reported as weighted accuracy.
@@ 5.1 STEL Performance
[…]
Table 1 reports style representation effectiveness (STEL) across training regimes Wegmann and Nguyen (2021). Across the board, iBERT models outperform their BERT counterparts, underscoring the utility of structured, interpretable pooling for capturing stylistic signals. Under SD-only training, iBERT-v3-1 achieves the highest STEL scores, yielding 38.3% (128 tokens) and 38.2% (512 tokens)—a substantial gain of +6.7 and +7.6 points over BERT. Close contenders include iBERT-v1 and iBERT-v3-10, both of which exhibit similarly strong performance, suggesting that soft, sense-weighted pooling is especially effective when guided by direct-style supervision.
@@
Even under joint supervision from WG+SD, where noisy author-level triplets may dilute sense-specific stylistic signal, iBERT variants maintain their edge. iBERT-v1 secures the best STEL at 128 tokens (38.6%), while iBERT-v3-1 leads at 512 tokens (36.4%), outperforming BERT by approximately 6 points in both cases.
Finally, in the WG-only setting, where labels reflect author identity rather than explicit styles, all models see degraded performance. Nevertheless, iBERT-v3 and iBERT-v1 is at par with BERT, signaling that even without style supervision, iBERT retains competitive structure-aware generalizability. These results affirm that iBERT’s compositional inductive bias enables robustness without compromising direct-style effectiveness.
@@ 5.2 SoC Performance
In contrast, iBERT-v2 underperforms across settings (e.g., 88.7% under SD-only), often lagging behind both v1 and BERT. We attribute this to its top-1 sense pooling: by forcing each sequence to align with a single dominant sense, iBERT-v2 severely restricts the number of senses that can specialize for the same data-points observed. This architectural bottleneck reduces expressive capacity, especially for fine-grained polarity distinctions that require distributed compositionality.
Under WG-only training—where supervision is provided weakly via author labels rather than direct stylistic polarity—all models see a dramatic drop in SoC performance. BERT, for example, falls from 91.5% (SD-only) to just 27.5%, while iBERT variants drop even further (e.g., 8.0% for v1). Despite this collapse, BERT appears to outperform iBERT in this setting. We hypothesize this is due to content leakage: BERT implicitly entangles semantic and stylistic attributes, using topic or lexical cues as proxies for style. While this can yield fragile SoC signal in weakly supervised settings, it sacrifices interpretability and modularity. In contrast, iBERT’s explicit disentanglement fails to recover polarity distinctions without true stylistic supervision—a tradeoff aligned with its core design goals.
@@ 5.3 Authorship Verification (PAN)
WG training dataset is a mixed dataset, containing author identity supervision with intertwined semantic and stylistic signals. Unlike BERT, which entangles all available cues, iBERT is architecturally designed to tease apart these differentials in an interpretable fashion. This enables it to maintain competitive performance even when the supervision is noisy or not explicitly style-specific—while still excelling on clean, style-focused tasks like STEL and SoC.
At 128 tokens, all iBERT variants match or slightly outperform BERT, with iBERT-v3-10 achieving the highest AUC (59.87%) for WG. This is notable not because iBERT ignores semantic information, but because it selectively organizes it in disentangled form—generalizing well even when stylistic cues are weakly defined.
At 512 tokens, BERT slightly outperforms iBERT (61.92% vs. 61.07% for v3), with the largest margins on PAN13 and PAN15, that are characterized by semantic drift and cross-topic variation. This suggests that BERT benefits more from extended context, due to its architecture’s tight coupling of semantic and stylistic features. In contrast, iBERT’s modular structure—designed to isolate stylistic axes—does not directly exploit semantic continuity. Still, the gap stays slim, highlighting the semantic signals acquired in MLM stage remain accessible even in disentangled representations.
@@ 5.4 Interpretability Analysis of iBERT
To validate these groupings, Table 3 reports the change in cosine distance between opposing style centroids before and after removing each sense contribution in final representation. When ablating a style-aligned sense, the separability between style polarities drops substantially (e.g., $66\%$ for $\ell=3$), confirming that key stylistic information is isolated within that sense. In contrast, most non-target styles exhibit low or negligible change in distance, indicating minimal collateral disruption: a hallmark of localized specialization.
@@
5.5 Visualize Targeted Editing via Ablation
@@
5.6 Visualize Non-Target Style Ablation
5.7 Quantifying Locality of Sense-Editing
A core utility for controllable representations is locality: editing a target attribute should only affect that attribute while minimally disturbing unrelated dimensions. While Figure 2 illustrates targeted edits qualitatively, we now formalize this behavior.
For a given style $x$ (e.g., Metaphorical / Literal), we identify the sense $\ell^{*}$ with the highest probing activation and ablate it during encoding by zeroing the corresponding sense gain. Let $P_{x}$ denote the positive examples for style $x$, and $P_{x}^{\text{orig}}$, $P_{x}^{\text{edit}}$ be their embeddings before and after ablating $\ell^{*}$, respectively. We compute cosine distance between a set of embeddings $A$ and a centroid $\mu$ as:
$d(A,\mu)=\frac{1}{|A|}\sum_{a\in A}\left(1-\cos(a,\mu)\right).$
The target editability is then:
$\Delta_{x}=\frac{d(P_{x}^{\text{orig}},\mu_{x}^{-})-d(P_{x}^{\text{edit}},\mu_{x}^{-})}{d(P_{x}^{\text{orig}},\mu_{x}^{-})},$
where $\mu_{x}^{-}$ is the centroid of negative examples for style $x$.
To measure non-target editing (collateral effects), we evaluate how the edits for $x$ impact other stylistic properties $y\neq x$, using $\mu_{y}^{-}$, which is the centroid of non-target $y$’s negative examples:
$\Delta_{y}=\frac{d(P_{x}^{\text{orig}},\mu_{y}^{-})-d(P_{x}^{\text{edit}},\mu_{y}^{-})}{d(P_{x}^{\text{orig}},\mu_{y}^{-})}.$
We report the following aggregate locality metrics:
Avg. Other Shift $=\frac{1}{|\{y\neq x\}|}\sum_{y\neq x}\Delta_{y}$
A useful and disentangled iBERT sense should ideally yield a large, positive $\Delta_{x}$ while keeping Avg. Other Shift low, indicating minimal collateral influence on unrelated style attributes.
Table 3 shows that ablating a style-aligned sense yields large reductions in target separability (up to 65%), while inducing minimal average shift on non-target styles. This confirms that stylistic attributes are concentrated in specific sense dimensions rather than distributed diffusely across the embedding.
As an added control, we refer again to Figure 3, which shows that ablating a non-primary sense (i.e., not the top sense $\ell^{*}$) for a given style leaves its attribute structures intact. This supports the view that iBERT’s sense representations are modular, with sense edits affecting the intended attributes, rather than causing broad changes across the space.
(a) $\ell=0$ ablated that aligns with More/Less conjunctions
(b) $\ell=2$ ablated that aligns with Sarcasm / No sarcasm
Figure 3: All Upper Case/Proper Capitalization style ($\ell^{*}$=1) maintains separation on ablating other senses ($\ell$={0, 2}). Colors as in Fig. 2- Red: original, Blue: edited, Gray: negatives.
@@ 6 Discussion
Ablating a single sense (e.g., $\ell{=}1$ for emoji, or $\ell{=}4$ for nominalization) shifts the representation toward the negative style without disrupting unrelated styles. These transformations are visible both numerically (Table 3) and spatially (Fig. 2), confirming that iBERT learns axis-aligned subspaces that can be controlled with precision. This kind of intervention is infeasible in dense encoders like SBERT or BERT.
@@ 7 Conclusion
iBERT matches or outperforms dense baselines across stylistic and semantic tasks like STEL, SoC, and PAN, while enabling: (a) disentangled encoding of style and semantics via sense-probing, (b) precise edits (e.g., formality), and (c) alignment of emergent senses with sociolinguistic axes.
@@
Model and Code
We release the models, demonstration, and code for other researchers to use. These are made available at https://github.com/vishalanand/iBERT and https://iBERT.io .
Limitations
@@
iBERT is trained on public data, including synthetic edits (StyleSynth) and large-scale web corpora. As with all pretrained models, underlying biases in this data may influence the dimensions that emerge. We encourage careful evaluation before deployment in applications with societal or demographic impact.
@@ Appendix A Appendix
Figure 4: t-SNE projections of sentence embeddings for SynthSTEL. iBERT consistently separates contrastive style variants (e.g., misspelled sentences, emoji usage) better than BERT, showing clearer margins and lower entanglement. Blue points represent positive samples and gray crosses represent negative samples. Despite iBERT-v2 being the most underperforming iBERT variant it still matches BERT’s performance, and the separation of positive and negative styles is cleaner.
@@
Table 5: Effect of aggregation sharpness $\tau$ on iBERT, trained on StyleSynth triplets (128 token size).
@@
Table 6: [iBERT-v3-10 (128-tokens)] Probing activations for style classes across all senses ($\ell=0$ to 77). The Best column indicates the latent sense with the highest activation for each style. Style names in this table reflect the original labels from the StyleSynth dataset. In the main text, shorter style names are used for presentation clarity.